how do i write two different expressions to find the total cost of horsebackriding.
Suppose you pay $15 per hour to go horseback riding. You ride 2 more hours today and plan to ride 4 more hours this weekend.


Answer by stanbon(75887)

how do i write two different expressions to find the total cost of horsebackriding.
------

Suppose you pay $15 per hour to go horseback riding. You ride 2 more hours today and plan to ride 4 more hours this weekend.
---
Todays cost = 2*15 = $30
------
Total cost = 30 + 4*15
Total cost = 30 + 60 = $90
==============================
cheers
Stan H.